OpenStudy (anonymous):

(sinθ)/(1-cosθ)*(1+cosθ)/(1+cosθ)

OpenStudy (richyw):

what are you asking? it equals \[\frac{\sin\theta}{1-\cos\theta}\]

OpenStudy (anonymous):

I'm asking to multiply it.

OpenStudy (richyw):

you are multiplying it by one though.

OpenStudy (anonymous):

oh ok thank you

OpenStudy (richyw):

assuming you formatted that correctly.

OpenStudy (richyw):

that's also \[\cot\left(\frac{\theta}{2}\right)\]